You are in a car moving at 90 km/h, and you see someone at the side of the highway. You use your horn to warn the person. The frequency of the horn is 440 Hz, and the air temperature is 0 °C. Calculate the frequency of the horn the person hears as you approach and as you pass.
Answer #1

Given

speed of car is vs = 90 kmph = 25 m/s

Speed of observer is zero,vo = 0 m/s

Frequency of the horn f = 440 Hz

Speed of sound v = 331.5  m/s

From Doppler effect , when the approaches the person, the frequency ,observed is

f'=f((v-vo)/(v-vs))

Substituting the values

f'=440((331.5-0)/(331.5-25))

f'= 475.89 Hz

When the car moves away from the person,the frequency observed by the person is

f'= f ((v-vo)/(v+vo))

Substituting the values

f'= 440((331.5-0)/(331.5+25)) Hz

f'= 409.14 Hz